OpenStudy (anonymous):

The earth is about 150 million kilometers from the sun. The orbit around the sun is not a perfect circle, but is approximately a circle. If the orbit of the earth around the sun were circular, what would be the equation if the origin of the coordinate system were placed at the sun?

OpenStudy (anonymous):

The equation for a circle centered at the origin with radius r is: \[x^2 + y^2 = r^2\] Since the radius is 150 million kilometers, we can plug that in for r and we'll get: \[x^2 + y^2 = (150 million)^2\] \[x^2 + y^2 = 2.25 \times 10^{16}\] Assuming you are graphing with kilometers as your units.

OpenStudy (anonymous):

thankk youu

OpenStudy (anonymous):

You're welcome, glad I could help.

OpenStudy (anonymous):

Here is your medal! You were right and I was thinking that it would need to end with a fraction. Thank you, lets work another on I will post it now.